# Quantum nonlinear effects in the number-conserving analogue gravity of Bose-Einstein condensates

### v1

@@ Abstract
We consider the number-conserving unitary dynamics of a Bose-Einstein condensates at absolute zero, and argue that an analogue gravity model in such a setting must take into account the backreaction of quasiparticle excitations onto the condensate background. This in turn requires that one expands to second order in perturbations, and takes the nonlinearity of the theory into account. It is shown that this leads to significant modifications of the standard linearized analogue gravity paradigm à la Unruh. In particular, to obtain a fully Lorentz-covariant equation in curved spacetime for second-order perturbations, we demonstrate that it is necessary to introduce, to leading order in powers of the formal expansion parameter $N^{-1/2}$ (where $N$ is total particle number), a quantum-fluctuation-renormalized spacetime metric which substantially differs from the Unruh metric and, to subleading order, two emergent vector fields. Both the renormalized metric and the vector fields then keep track of the effect of the backreaction of the quasiparticles onto the condensate up to the order in powers of $N^{-1/2}$ considered. Finally, we apply our formalism to an analogue-cosmological Friedmann-Lemaître-Robertson-Walker metric and establish its renormalized form due to quantum many-body backreaction.
@@ I Introduction
One crucial aspect of probing quantized fields in an arbitrary curved spacetime that is often neglected is the possible effect of the backreaction of the fluctuating field on the background itself. The relative magnitude of backreaction, say in terms of the contribution to the semiclassical energy-momentum tensor, must be much smaller than that of the gravitating matter field for a clear separation between what constitutes “background” and “fluctuation” to be maintained. To the lowest order, as in the original prediction of the Hawking effect, backreaction is indeed conventionally neglected [24], or at most a toy model for backreaction is proposed [25, 26]. However, a completely self-consistent treatment of any quantum field theory of a many-body system (or, similarly, of any candidate for an emergent unitary quantum field theory of gravity), should accommodate backreaction due to quantum-mechanical fluctuations created by its dynamical evolution. In this regard, analogue gravity models based on BECs represent an ideal testbed to explore the possible role of quantum backreaction, as these systems afford a microscopically fully controlled description in terms of Bogoliubov theory [27, 28].
@@
It is of major importance in what follows that for a proper treatment of backreaction problems in quantum systems such as BECs, as was first pointed out in [29], it is necessary to go beyond the usual symmetry-breaking approach violating particle number conservation, and use a manifestly number-conserving description of the quantum many-body system [30, 31, 32, 33]. To this end, in [29] the backreaction problem was defined by employing an expansion of the field operator in powers of $N^{1/2}$, where $N$ is the total number of particles in the system. Then the backreaction of the quantum fluctuations can be elegantly treated as the subleading correction to the leading order in $N^{1/2}$ Gross-Pitaevskiǐequation(̃GPE) [equivalent to a nonrelativistic variant of $\phi^{4}$ theory], incorporated by the modified coupled dynamical equation of the condensate and the non-condensate part. A particular lesson to be drawn from the analysis of [29] is that in a fully number-conserving description of the backreaction problem, the usual assumption of taking the expectation value of the quantized energy-momentum tensor, $\hat{T}_{\mu\nu}$, in a suitable quantum state, cf. Ref. [34], does not necessarily reproduce the correct, that is observable physical backreaction force. This approach was further explored in [35], where the solution for backreaction as a Cauchy problem was formulated, and solved for a dynamical BEC configuration. Importantly, it was shown in [35, 36] that there can exist a nonzero flux of condensate particles even in the absence of any particle flux due to quantum depletion, solely due to backreaction. Backreaction for a solitonic solution of the nonlinear Schrödinger equation using the number-conserving Bogoliubov theory was also studied most recently in [37].
@@
In this work, our aim is to build an analogue gravity model based on the fully number-conserving approach incorporating the coupled dynamics (particle exchange) of the condensate and the depleted cloud. As we will explain in the sequel, our work is based on two key features that emerge from a consistent treatment of the backreaction of the depleted cloud on the background [29, 38, 39, 35]. Firstly, due to number conservation, there is a nonvanishing particle flux from the condensate to the depletion. As a result, in the hydrodynamical approach, the continuity equation of the fluid variables (density and velocity potential) contains drain (source) terms. And secondly, the background to build an analogue gravity description now has a subleading contribution that represents a correction to the usual GPE. Consequently, as we argue, the standard approach of using strictly linear perturbation fields (see for example Visser’s pedagogical paper [3]) is not consistent with the order (in the perturbation expansion in $N$) of the field operator we work with, ${\mathcal{O}}(N^{-1/2})$, which is, as we show, the minimal expansion order necessary for number conservation to be maintained. This requires us to go beyond the linear perturbation paradigm. On the classical level, this was developed in Ref. [40], where it was shown how to extract an effective metric description from a generic nonlinear perturbation, at the cost of modifying the definition of background, also cf. Ref. [41].
@@ II.1 Governing field equations
$i\partial_{t}\Psi=\left(-\frac{1}{2m}\nabla^{2}+V+U|\Psi|^{2}\right)\Psi,$ (1)
Here, $V$ is the external (trapping) potential, and $U$ is the interaction coupling from $s$-wave scattering. As can be easily recognized, the theory described by (1) is invariant under global $U(1)$ transformations, which ensures the conservation of the total particle number of the system $N$ according to the Noether theorem. The conservation law $\partial_{t}\rho+\nabla\cdot{\bm{J}}=0$ thus holds, where
$\rho=|\Psi|^{2},\qquad{\bm{J}}=\frac{1}{2im}(\Psi^{*}\nabla\Psi-\Psi\nabla\Psi^{*})$ (2)
@@
The variant of number-conserving theory we work with relies on the following expansion in powers of $N^{1/2}$
$\Psi=\phi_{\rm 0}+\chi+\phi_{2}+\mathcal{O}(N^{-3/2}).$ (3)
@@
$i\frac{\partial}{\partial t}\phi_{0}=\left[-\frac{1}{2m}\nabla^{2}+V+U|\phi_{0}|^{2}\right]\phi_{0}.$ (5)
and the modified Bogoliubov–de Gennes (BdG) equation for the number-conserving fluctuation field $\chi$
$i\frac{\partial}{\partial t}\chi=\left(-\frac{1}{2m}\nabla^{2}+V+2U\rho_{0}\right)\chi+U\phi_{\rm 0}^{2}\chi^{*},$ (6)
@@
$i\partial_{t}\phi_{c}=\Big{(}-\frac{1}{2m}\nabla^{2}+V+U|\phi_{\rm c}|^{2}+2U|\chi|^{2}\Big{)}\phi_{c}+U\chi^{2}\phi_{c}^{*}.$ (8)
@@
To formulate the quantization of the fluctuation on top of the classical background, we have to promote the fluctuation field $\chi$ and its Hermitian conjugate to be quantum operators satisfying the standard Bose commutation relations. We will also assume the vacuum state to be defined by the quasiparticle vacuum $\langle\hat{\chi}\rangle\coloneqq 0$. Then the equation governing the dynamics of the corrected mean field is given by the expectation value of (8) in this state,
$i\partial_{t}\phi_{c}=\left[-\frac{1}{2m}\nabla^{2}+V+U|\phi_{\rm c}|^{2}+2U\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle\right]\phi_{c}+U\langle\hat{\chi}^{2}\rangle\phi_{c}^{*}.$ (9)
We will not delve into the details of defining the quasiparticle operator but will rather take the (pragmatic) viewpoint that field quantization can be performed at least for cases where the quasiparticle vacuum is well defined, such as in the simplest, that is noninteracting case. We therefore assume that we can evaluate terms such as the anomalous density appearing in the modified GPE by solving the $\mathcal{O}(N^{0})$ BdG equation, as was done for example in [35, 36], as an explicit functional of the leading order GPE orbital $\phi_{0}$, along with the associated dependence on the spacetime coordinates $x,t$, see the Appendix for details on how to carry this out for the setup studied in [35, 36].
@@ II.2 Quantum backreaction
We summarize here the formulation of the quantum backreaction problem in a fully number-conserving formalism, taking the ${\mathcal{O}}(N^{-1/2})$ correction to the condensate background [29] into consideration. The full system satisfies a conservation equation, $\partial_{t}\rho+\nabla\cdot{\bm{J}}=0$, where the average density is given as
@@
${\bm{J}}:=\langle\hat{\bm{J}}\rangle=\frac{1}{2im}\left\langle\hat{\Psi}^{\dagger}\nabla\hat{\Psi}-\nabla\hat{\Psi}^{\dagger}\hat{\Psi}\right\rangle.$ (11)
Utilizing the expansion of the full field operator (3) and evaluating the total density and current operator (2) in the quasiparticle vacuum, we can write the density and current equations as
$\rho$ $=$ $|\phi_{c}(\mathbf{x})|^{2}+\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle+{\mathcal{O}}(N^{-1/2}),$ (12)
${\bm{J}}$ $=$ $\frac{1}{2im}\Big{(}\phi^{*}_{c}\nabla\phi_{c}(\mathbf{x})-\nabla\phi^{*}_{c}\Big{)}$ (13)
$+$ $\frac{1}{2im}\left\langle\hat{\chi}^{\dagger}\nabla\hat{\chi}-\nabla\hat{\chi}^{\dagger}\hat{\chi}\right\rangle+{\mathcal{O}}(N^{-1/2}).$
We have ordered the various terms by their order in $N^{1/2}$ For a proper description of the condensate correction due to quantum fluctuations it is useful to separate the contribution of the condensate and the noncondensate part in the full density and current as follows
@@
${\bm{J}}_{\hat{\chi}}=\frac{1}{2im}\left\langle\hat{\chi}^{\dagger}\nabla\hat{\chi}-\nabla\hat{\chi}^{\dagger}\hat{\chi}\right\rangle,$ (15)
@@
Neither condensate nor depletion are separately conserved, if even only up to the order ${\mathcal{O}}(N^{0})$ we consider. In particular, we have for condensate and depleted clouds, respectively [35, 36],
@@
We conclude that the theory is indeed conserving up to order ${\mathcal{O}}(N^{0})$ for density and current, but the separate conservation equations for condensate and excitations acquire higher-order drain (source) terms in a number-conserving approach such as ours.
To summarize, we have derived that the number of particles in neither condensate nor in quantum excitations are separately conserved. It is thus only meaningful to consider their coupled dynamics to understand the backreaction of the quantum fluctuations onto the condensate.
III Analogue gravity in the number-conserving formalism
@@
$\partial_{t}\rho_{c}+\frac{1}{m}\nabla\cdot\left(\rho_{c}\nabla\theta_{c}\right)=f_{1}(\rho_{c},\theta_{c},{\bm{x}},t),$ (19)
$\begin{split}\partial_{t}\theta_{c}+U\rho_{c}+\frac{1}{2m}\left(\nabla\theta_{c}\right)\cdot\left(\nabla\theta_{c}\right)+f_{3}(\rho_{c},\theta_{c},{\bm{x}},t)\\ +V({\bm{x}},t)=0.\end{split}$ (20)
@@
which are related to real and complex parts of the “anomalous” density. We have omitted the “quantum pressure” term $-h(\rho_{c})=-\frac{1}{2m}\rho_{c}^{-1/2}\left(\nabla^{2}\rho_{c}^{1/2}\right)$ on the left-hand side of the Bernoulli Eq. (20); we will generally neglect density gradients in the various relevant background configurations (see below) in what follows. Including these gradients would lead us to the concept of (nonlocal) rainbow metrics [45]. The indices 1 and 3 on $f_{1}$ and $f_{3}$ indicate their respective orders ${\cal O}(N^{0})$ and ${\cal O}(N^{-1})$; see also the note after Eq. (24) below.
In the remainder of the paper, we will work with (19) and (20), which are essentially the standard fluid dynamical quasi-nonlinear equations, based on which we will explore our construction of analogue gravity in the second order nonlinear response regime. It is important to point out that the two functions $f_{1}$ and $f_{3}$ which contain the deviation from the usual hydrodynamic description based on the usual leading order GPE are of the order $\mathcal{O}(N^{0})$ and $\mathcal{O}(N^{-1})$, respectively. Thus they are subdominant contributions to Eqs. (19 and 20). They however encapsulate the backreaction we are interested in, and thus we fully keep $f_{1},f_{3}$, assuming them to be general functions of $\rho_{c}$, and $\theta_{c}$. Note that $f_{1}<0$ corresponds to the loss of particles from the condensate.

### v2

@@ Abstract
We consider the unitary quantum dynamics of a Bose-Einstein condensates at absolute zero, and demonstrate that an analogue gravity model in such a setting must take into account the backreaction of quasiparticle excitations onto the condensate background. This in turn requires that one expands to second order in perturbations, and takes the nonlinearity of the theory into account. It is shown that this leads to significant modifications of the standard linearized analogue gravity paradigm à la Unruh. In particular, to obtain a fully Lorentz-covariant equation in curved spacetime for second-order perturbations, we demonstrate that it is necessary to introduce, to leading order in powers of the formal expansion parameter $N^{-1/2}$ (where $N$ is total particle number), a quantum-fluctuation-renormalized spacetime metric which substantially differs from the Unruh metric and, to subleading order, two emergent vector fields. Both the renormalized metric and the vector fields then keep track of the backreaction of the quasiparticles onto the condensate up to the order in powers of $N^{-1/2}$ considered. Finally, we apply our formalism to an analogue-cosmological Friedmann-Lemaître-Robertson-Walker metric and establish its renormalized form due to the quantum many-body backreaction exerted by the excitation cloud.
@@ I Introduction
One crucial aspect of probing quantized fields in an arbitrary curved spacetime that is often neglected is the possible effect of the backreaction of the fluctuating field on the background itself. The relative magnitude of backreaction, say in terms of the contribution to the semiclassical energy-momentum tensor, must be much smaller than that of the gravitating matter field for a clear separation between what constitutes “background” and “fluctuation” to be maintained. To the lowest order, as in the original prediction of the Hawking effect, backreaction is indeed conventionally neglected [24], or at most a toy model for backreaction is proposed [25, 26]. However, a completely self-consistent treatment of any quantum field theory of a many-body system (or, similarly, of any candidate for an emergent unitary quantum field theory of gravity), should accommodate backreaction due to quantum-mechanical fluctuations created by its dynamical evolution. In this regard, analogue gravity models based on BECs represent an ideal testbed to explore the possible role of quantum backreaction, as these systems afford a fully controlled description in terms of Bogoliubov theory [27, 28, 29].
@@
It is of major importance in what follows that for a proper treatment of backreaction problems in quantum systems such as BECs, as was first pointed out in [30], it is necessary to maintain particle-number conservation, cf. Refs. [31, 32, 33, 34], which are obtaining variants of number-conserving formulations of Bogoliubov theory which go beyond the usual symmetry-breaking, traditional Bogoliubov approach. To this end, in [30] the backreaction problem was defined by employing an expansion of the field operator in powers of $N^{1/2}$, where $N$ is the total number of particles (here either atoms or molecules) in the system. Then the backreaction of the quantum fluctuations can be elegantly treated as the subleading correction to the leading order in $N^{1/2}$ Gross-Pitaevskiǐ equation (GPE) [equivalent to a nonrelativistic variant of $\phi^{4}$ theory], incorporated by the modified coupled dynamical equation of the condensate and the non-condensate part. A particular lesson to be drawn from the analysis of [30] is that in a fully number-conserving description of the backreaction problem, the usual assumption of taking the expectation value of the quantized energy-momentum tensor, $\hat{T}_{\mu\nu}$, in a suitable quantum state, cf. Ref. [35], does not necessarily reproduce the correct, that is observable physical backreaction force.
The consequences of using a number-conserving approach were further explored in [36], where the solution for backreaction as a Cauchy problem was formulated, and solved for a dynamical BEC configuration. Importantly, it was shown in [36, 37] that there can exist a nonzero flux of condensate particles even in the absence of any particle flux due to quantum depletion, solely due to backreaction. Backreaction for a solitonic solution of the nonlinear Schrödinger equation was also studied most recently in [38].
@@
In the present study, our aim is to build an analogue gravity model based on the number-conserving approach of [38], which incorporates the coupled dynamics (particle exchange) of the condensate and the depleted cloud. As we will explain in the sequel, the following is based on two key features that emerge from a consistent treatment of the backreaction of the depleted cloud onto the background [30, 39, 40, 36]. First of all, due to number conservation, there is a nonvanishing particle flux from the condensate to the depletion part. As a result, the continuity equations for density and current of both condensate and depletion contain drain (source) terms. In addition, the background to build an analogue gravity description now has a subleading contribution that represents a correction to the usual GPE. Consequently, as we argue, the standard approach of using strictly linear perturbation fields (see for example Visser’s pedagogical paper [3]) is not consistent with the order (in the perturbation expansion in $N$) of the field operator we will work with, ${\mathcal{O}}(N^{-1/2})$, which will be demonstrated to be the minimal expansion order necessary for number conservation to be maintained. We are thus required to go beyond the linear perturbation paradigm of Unruh. On the classical level, the latter was developed in Ref. [41], where it was shown how to extract an effective metric description from a generic nonlinear perturbation, at the cost of modifying the definition of background, also cf. Ref. [42].
@@ II.1 Governing field equations
$i\partial_{t}\Psi=\left(-\frac{1}{2}\nabla^{2}+V+U|\Psi|^{2}\right)\Psi,$ (1)
Here, $V$ is the external (trapping) potential, and $U>0$ is the interaction coupling from $s$-wave scattering between the gas particles. The theory described by (1) is invariant under global $U(1)$ transformations, which ensures the conservation of the total particle number $N$ according to the Noether theorem. The conservation law $\partial_{t}\rho+\nabla\cdot{\bm{J}}=0$ thus holds, where
$\rho=|\Psi|^{2},\qquad{\bm{J}}=\frac{1}{2i}(\Psi^{*}\nabla\Psi-\Psi\nabla\Psi^{*})$ (2)
@@
The variant of number-conserving theory we work with is based on the following formal expansion of the classical field in powers of $N^{1/2}$ [^1] Note there is a typo in Eq. (2) of Ref. [36], where ${\cal O}(N^{-3/2})$ should be replaced by $\mathcal{O}(N^{-1})$
$\Psi=\phi_{\rm 0}+\chi+\phi_{2}+\mathcal{O}(N^{-1}).$ (3)
@@
$i\frac{\partial}{\partial t}\phi_{0}=\left[-\frac{1}{2}\nabla^{2}+V+U|\phi_{0}|^{2}\right]\phi_{0}.$ (5)
and the modified Bogoliubov–de Gennes (BdG) equation for the fluctuation field $\chi$
$i\frac{\partial}{\partial t}\chi=\left(-\frac{1}{2}\nabla^{2}+V+2U\rho_{0}\right)\chi+U\phi_{\rm 0}^{2}\chi^{*},$ (6)
@@
$i\partial_{t}\phi_{c}=\Big{(}-\frac{1}{2}\nabla^{2}+V+U|\phi_{\rm c}|^{2}+2U|\chi|^{2}\Big{)}\phi_{c}+U\chi^{2}\phi_{c}^{*}.$ (8)
@@
To formulate the quantization of the fluctuation on top of the classical background constituted by $\phi_{c}$, we promote the fluctuation field $\chi$ and its Hermitian conjugate to be the fundamental quantum operators satisfying the standard bosonic commutation relations
$\left[\hat{\chi}({\bm{x}},t),\hat{\chi}^{\dagger}({\bm{x}}^{\prime},t)\right]=\delta({\bm{x}}-{\bm{x}}^{\prime}),$
$\left[\hat{\chi}({\bm{x}},t),\hat{\chi}({\bm{x}}^{\prime},t)\right]=\left[\hat{\chi}^{\dagger}({\bm{x}},t),\hat{\chi}^{\dagger}({\bm{x}}^{\prime},t)\right]=0.$ (9)
We will then assume that the instantaneous quantum state under consideration fulfills
$\langle\hat{\chi}({\bm{x}},t))\rangle\coloneqq 0\,\,\,\forall\,{\bm{x}},t\geq 0.$ (10)
This quantization choice, that is to regard $\hat{\chi}$ as the fundamental bosonic quantum field instead of $\hat{\Psi}$, does not maintain number conservation on the level of the full field operator of the bosonic particles: Due to $\langle\hat{\Psi}\rangle=\phi_{c}$, where however now $\phi_{c}$ is the backreaction-renormalized condensate mode in Eq. (7), the U(1) symmetry is still broken on the quantum level, as in the original Bogoliubov theory [27]. This choice, to promote $\chi$ to be the “fundamental” quantum field, however yields a set of equations amenable to a concrete solution the backreaction problem  [^2] We note in this connection that we were unable to locate an explicit backreaction calculation following the formally rigorous and in this regard very appealing approach of Ref. [33], which maintains $\langle\hat{\Psi}\rangle=0$ for any fixed $N$ state.. Furthermore, and most importantly in our context, it has been shown in [36] that with this vacuum choice for the quantized field $\hat{\chi}$, number conservation for condensate plus excitations is maintained, see Eqs. (18) and (19) below, and up to the order in $N^{-1/2}$ considered. The initial state for the Heisenberg picture evolution is then chosen as the noninteracting $U=0$ which has zero depletion $\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle=0\,\,(t=0)$ and $\phi_{2}=0$ (zero backreaction) [36] and can be set up in experiment in a well controlled manner. At $t=0$, we quench to a finite value of $U$ and follow the subsequent evolution.
The equation governing the dynamics of the corrected mean field is then given by the expectation value of (8)
$i\partial_{t}\phi_{c}=\left[-\frac{1}{2}\nabla^{2}+V+U|\phi_{\rm c}|^{2}+2U\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle\right]\phi_{c}+U\langle\hat{\chi}^{2}\rangle\phi_{c}^{*}.$ (11)
We will not delve into the details of defining the quasiparticle operator but will rather take the (pragmatic) viewpoint that field quantization can be performed at least for cases where the quasiparticle vacuum is well defined, such as in the simplest, that is noninteracting case. We therefore assume that we can evaluate terms such as the anomalous density appearing in the modified GPE by solving the $\mathcal{O}(N^{0})$ BdG equation as an explicit functional of the leading order GPE orbital $\phi_{0}$, along with the associated dependence on the spacetime coordinates $x,t$: see Appendix for a detailed calculation with the setup of Refs. [36, 37].
@@ II.2 Quantum backreaction
We summarize here the formulation of the quantum backreaction problem in the formalism of Ref. [36], by taking the ${\mathcal{O}}(N^{-1/2})$ correction to the condensate background into consideration. The full system satisfies a conservation equation, $\partial_{t}\rho+\nabla\cdot{\bm{J}}=0$, where the average density is given as
@@
${\bm{J}}:=\langle\hat{\bm{J}}\rangle=\frac{1}{2i}\left\langle\hat{\Psi}^{\dagger}\nabla\hat{\Psi}-\nabla\hat{\Psi}^{\dagger}\hat{\Psi}\right\rangle.$ (13)
Utilizing the expansion of the full field operator (3) and evaluating the expressions above given the condition (10) we can write the density and current equations as
$\rho$ $=$ $|\phi_{c}({\bm{x}}|^{2}+\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle+{\mathcal{O}}(N^{-1/2}),$ (14)
${\bm{J}}$ $=$ $\frac{1}{2i}\Big{(}\phi^{*}_{c}\nabla\phi_{c}({\bm{x}}-\nabla\phi^{*}_{c}\Big{)}$ (15)
$+$ $\frac{1}{2i}\left\langle\hat{\chi}^{\dagger}\nabla\hat{\chi}-\nabla\hat{\chi}^{\dagger}\hat{\chi}\right\rangle+{\mathcal{O}}(N^{-1/2}).$
We have ordered the various terms by their order in $N^{1/2}$. For a proper description of the condensate correction due to quantum fluctuations it is useful to separate the contribution of the condensate and the noncondensate part in the full density and current as follows
@@
${\bm{J}}_{\hat{\chi}}=\frac{1}{2i}\left\langle\hat{\chi}^{\dagger}\nabla\hat{\chi}-\nabla\hat{\chi}^{\dagger}\hat{\chi}\right\rangle,$ (17)
@@
Neither condensate nor depletion are separately conserved, even if only up to the order ${\mathcal{O}}(N^{0})$ we consider. In particular, we have for condensate and depletion cloud, respectively [36, 37],
@@
We conclude that the theory is indeed conserving up to order ${\mathcal{O}}(N^{0})$ for density and current. This defines the extent to which our approach is number-conserving, as required. The number of particles in condensate and in the depletion cloud of quantum excitations are not separately conserved but only their respective sums, and thus the total density and current. The separate conservation equations for condensate and excitations acquire drain (source) terms which stem from backreaction.
III Analogue gravity in a number-conserving formalism
@@
$\partial_{t}\rho_{c}+\nabla\cdot\left(\rho_{c}\nabla\theta_{c}\right)=f_{1}(\rho_{c},\theta_{c},{\bm{x}},t),$ (21)
$\begin{split}\partial_{t}\theta_{c}+U\rho_{c}+\frac{1}{2}\left(\nabla\theta_{c}\right)\cdot\left(\nabla\theta_{c}\right)+f_{3}(\rho_{c},\theta_{c},{\bm{x}},t)\\ +V({\bm{x}},t)=0.\end{split}$ (22)
@@
which are related to real and complex parts of the “anomalous” density. We have omitted the “quantum pressure” term $-h(\rho_{c})=-\frac{1}{2}\rho_{c}^{-1/2}\left(\nabla^{2}\rho_{c}^{1/2}\right)$ on the left-hand side of the Bernoulli Eq. (22); we will generally neglect density gradients in the background(s) in what follows. Including these gradients would lead us to the concept of (nonlocal) rainbow metrics [48]. The indices 1 and 3 on $f_{1}$ and $f_{3}$ indicate their respective orders ${\cal O}(N^{0})$ and ${\cal O}(N^{-1})$; see also the note after Eq. (26) below.
In the remainder of the paper, we will work with (21) and (22), which are essentially the standard fluid dynamical quasi-nonlinear equations, based on which we will explore our construction of analogue gravity in the second order nonlinear response regime. It is important to point out that the two functions $f_{1}$ and $f_{3}$ which contain the deviation from the usual hydrodynamic description based on the usual leading order GPE are of the order $\mathcal{O}(N^{0})$ and $\mathcal{O}(N^{-1})$, respectively. Thus they are subdominant contributions to Eqs. (21 and 22). They however encapsulate the backreaction we are interested in, and thus we fully keep $f_{1},f_{3}$, assuming them to be general functions of $\rho_{c}$, and $\theta_{c}$. Note that $f_{1}<0$ corresponds to the loss of particles from the condensate.